OpenStudy (anonymous):

Suppose that 320 feet of fencing are available to enclose a rectangular field and that one side of the field must be given double fencing. What are the dimenstions of the rectangle that will yield the largest possible area?

OpenStudy (anonymous):

if l = length and w = width then 3l + 2w = 320 2w = 320 - 3l w = 160 - 1.5l area A = w l = 160l - 1.5l^3 differentiate dA/ dl = 160 - 3l for max/min area da/dl = 0 160-3l = 0 l = 160/3 second derivative is -3 - negative so its a maxm maximum area has length 53.33 and width = 160 - 1.5 * 160/3 = 80 dimensions are 80 * 53.33 feet

OpenStudy (anonymous):

1) P = 2L + 3 W = 320 2) A = L * W Solve for L or W 160 - 3/2 W = L Substitute into equation 2 A(W) = 160W - 3/2 W^2 Maximum/Min will be found by the first derivative setting equal zero A'(W) = 160 - 3W = 0 W = 160/3 L = 80 Dimensions would be 80, 160/3

OpenStudy (anonymous):

pls reverse w and l

OpenStudy (anonymous):

Thanks guys. Only question is how you came up with the 3w and 2l in the beginning?